alibaby:

What is the value of the expression 22 + 42 ÷ 22? 3 6 8 16

alibaby:

@Logic007 @Aeon

Logic007:

\(\frac{2}{2}+\frac{4}{2}\div\frac{2}{2}\) like this? @alibaby